OpenStudy (anonymous):

Calculus III: find the center of mass of a 3-4-6 triangle by use of a double integral

OpenStudy (anonymous):

first we need to come up with functions

OpenStudy (anonymous):

alright. How do we go about that?

OpenStudy (anonymous):

are you sure it is not 3-4-5 triangle ?

OpenStudy (anonymous):

no, it's a 3-4-6

OpenStudy (anonymous):

I emailed him and he said something about law of cosines, b/c it's not a right triangle

OpenStudy (anonymous):

*emailed teacher

OpenStudy (anonymous):

alright, your professor sucks

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

alright, let's find one angle using law of cosines http://www.mathsisfun.com/algebra/trig-cosine-law.html 3^2=4^2+6^2-2(4)(6) cos(C) 9=16+36-48 cos(C) 43=48 cos(C) 43/48 = Cos(C) C=0.460493